# Double Robustness of Local Projections

### v1

@@ 3.3 Asymptotic coverage
See Section C.2. ∎
We again see that LP robustly controls coverage, while the VAR confidence interval generically has coverage converging to zero for $\zeta\in(1/4,1/2)$, and strictly below the nominal level $1-a$ for $\zeta=1/2$.
Role of the lag length.
A simple way to remove the asymptotic bias of the VAR estimator is to control for sufficiently many lags, since in this case the estimator is asymptotically equivalent with the LP estimator. The larger the horizon of interest, the more lags are required for bias reduction.
@@
See Section C.3. ∎
See Plagborg-Møller and Wolf (2021) and Xu (2023) for related results in models without explicit MA misspecification.
Hausman test.
To interpret the magnitude of the local misspecification, it is helpful to consider a Hausman (1978) test of correct specification of the VAR model that compares the VAR and LP impulse response estimates. This test rejects for large values of $\sqrt{T}|\hat{\beta}_{h}-\hat{\delta}_{h}|/\sqrt{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})}$. A test of this kind was proposed by Stock and Watson (2018) in the context of testing for invertibility.
Proposition 3.3.
where $b_{h}$ and $r(\cdot,\cdot)$ were defined in 3.2.
Proof.
@@
As claimed previously, the Hausman test is consistent against MA misspecification of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$, except in the knife-edge case $b_{h}=0$. When $\zeta=1/2$ and $b_{h}\neq 0$, the asymptotic rejection probability is strictly between the significance level $a$ and 1. In the next section we shall use the Hausman test to quantify the difficulty of detecting especially pernicious types of model misspecification.
@@ 4 Some unpleasant VARithmetic
We now argue that our theoretical results on VAR fragility are likely to have serious implications for VAR inference in practice. Consistent with the idea that finite-order VARs can provide a good—but not perfect—approximation of reality, we impose an a priori constraint on the magnitude of the MA misspecification. We then show that, even if the amount of misspecification is small, the worst-case bias of the VAR estimator is small if, and only if, its variance is similar to that of LP. When the relative standard errors of the VAR and LP estimators is in an empirically relevant range, the VAR confidence interval exhibits severe coverage distortions for misspecification that is small, hard to rule out ex ante on economic theory grounds, and also difficult to detect ex post with model validation tests. While the VAR under-coverage can be fixed by using a larger bias-aware critical value, we show that the resulting confidence interval is usually wider than the LP interval.
@@
4.1 Worst-case bias
@@
Now suppose we are willing to impose a priori that the noise-to-signal ratio is at most $M^{2}/T$ for some constant $M\in(0,\infty)$. For small $M^{2}/T$, this roughly means that a fraction $M^{2}/T$ of the variance of the model’s error term is due to the misspecification. This corresponds to restricting the parameter space for $\alpha(L)$ to all absolutely summable lag polynomials that satisfy $\|\alpha(L)\|\leq M$. In the following we will consider the worst-case properties of the VAR estimator over this parameter space, treating the other (consistently estimable) parameters $(A,H,D)$ as fixed.
@@
See Section B.4. ∎
Under our bound $M^{2}/T$ on the noise-to-signal ratio, the worst-case (scaled) VAR bias is a simple function only of $M$ and of the relative asymptotic precision $\operatorname{aVar}(\hat{\beta}_{h})/\operatorname{aVar}(\hat{\delta}_{h})$ of the VAR estimator vs. LP. These two quantities are “sufficient statistics” for the worst-case bias regardless of the number $n$ of variables in the VAR, the lag length $p$, the specific VAR parameters $(A,H,D)$, and the horizon $h$. Hence, our subsequent analysis of the worst-case properties of VAR procedures depends only on $M$ and on the relative precision, allowing us to concisely present analytical results that cover a wide range of local-to-SVAR models without having to resort to simulations that inevitably only cover a finite number of DGPs.
Proposition 4.1 shows that there is no “free lunch” for VAR estimation: the worst-case VAR bias is small precisely when the VAR estimator has nearly the same variance as LP. While the worst-case bias can be reduced by increasing the VAR estimation lag length $p$, the proposition shows that this can only happen at the expense of increasing the variance. If we include so many lags that the worst-case bias is zero (cf. 3.3), then the VAR estimator must necessarily be asymptotically equivalent with LP.
4.2 Worst-case mean squared error
To benchmark our later results on the worst-case properties of VAR inference procedures, we will first very briefly show how the worst-case mean squared error (MSE) of the VAR estimator depends on the imposed bound on misspecification. Based on Propositions 3.1 and 3.2 as well as 3.1, we define the asymptotic MSE of the VAR and LP estimators as follows:
@@ 4.2 Worst-case mean squared error
See Section C.4. ∎
In words, the worst-case MSE regret of VAR relative to LP is proportional to the variance reduction of VAR relative to LP, with proportionality constant $M^{2}-1$. If $M>1$ (corresponding to a noise-to-signal ratio greater than $1/T$), the worst-case MSE of VAR thus strictly exceeds the MSE of LP.
We can also ask what is the minimax optimal way to average LP and VAR estimates.
@@ 4.3 Worst-case coverage
We now turn to our main area of interest: the worst-case coverage of the conventional VAR confidence interval under small amounts of misspecification. We also characterize the least favorable type of misspecification and argue that it is difficult to rule out.
3.2 and 4.1 immediately imply an expression for the worst-case asymptotic coverage of the conventional VAR confidence interval.
@@
This is an immediate consequence of 3.2, 3.3 and 4.1. ∎
Figure 4.1 plots this worst-case coverage probability for different values of $M$, and given significance level $a=10\%$. We see that even for $M=1$ (corresponding to a noise-to-signal ratio of $1/T$), the worst-case coverage probability is below 48% whenever the asymptotic standard deviation of the VAR estimator is less than half that of LP, a standard error ratio commonly encountered in empirical work.
Figure 4.1: Worst-case asymptotic coverage probability of the conventional 90% VAR confidence interval. Horizontal axis: relative asymptotic standard deviation of LP vs. VAR. Different lines: different bounds $M$ on $\|\alpha(L)\|$. The solid horizontal line marks the nominal coverage probability $1-a=90\%$.
We conclude that, even for small amounts of VAR misspecification (as measured by the noise-to-signal ratio) and for relative estimator precisions of typical magnitudes, the scope for VAR undercoverage is material. This potential undercoverage may not be so concerning if the worst-case misspecification can be ruled out on economic theory grounds, or if it is easily detectable statistically. We now argue that neither appears to be the case.
@@
The shape and magnitude of the least favorable misspecification is difficult to rule out generally based on economic theory. The least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha_{\ell,h,M}^{\dagger}L^{\ell}$ for coverage is the same as the least favorable one for bias (i.e., the $\alpha(L)$ that achieves the maximum in Proposition 4.1). Since $\operatorname{aBias}(\hat{\delta}_{h})$ is linear in $\alpha(L)$, the least favorable choice given the constraint $\|\alpha(L)\|\leq M$ follows easily from the Cauchy-Schwarz inequality (see the proof of Proposition 4.1):
$\alpha_{\ell,h,M}^{\dagger}\propto D^{1/2}H^{\prime}\Psi_{h}^{\prime}S^{-1}A^{\ell-1}HD^{1/2}-\mathbbm{1}(\ell\leq h)\sigma_{j^{*}}^{-1}D^{1/2}H^{\prime}(A^{\prime})^{h-\ell}e_{i^{*},n}e_{j^{*},m}^{\prime},\quad\ell=1,2,\dots,$
@@
We note two main properties of the worst-case $\alpha_{\ell,h,M}^{\dagger}$. First, the magnitude of the MA coefficients decays exponentially as $\ell\to\infty$. In other words, not only is the overall magnitude of the least favorable model misspecification small (as imposed in the noise-to-signal bound), the MA coefficients at long lags are particularly small. Second, numerical examples shown in Section A.1 suggest that the MA coefficients tend to be largest in magnitude at horizon $h$, displaying either a hump-shaped pattern as a function of $\ell$—consistent with economic theories of adjustment costs or learning—or a single zig-zag pattern—consistent with theories of overshooting or lumpy adjustment. In fact, as we will show formally in Section 5.2, simple canonical models like an ARMA(1,1) map into misspecification polynomials $\alpha(L)$ that look extremely similar to the worst-case misspecification $\alpha^{\dagger}(L)$. We thus view MA dynamics of the worst-case form as empirically and theoretically realistic types of misspecification.
@@
Rather than committing a priori to a parameter space for $\alpha(L)$ through choice of $M$, we can also ask a different question: over all possible types and magnitudes of misspecification, what is the worst-case probability that the conventional VAR confidence interval fails to cover the true impulse response, yet we fail to reject correct specification of the VAR model?
@@
See Section C.6. ∎
Figure 4.2 plots this worst-case probability for significance level $a=10\%$, which as we have seen depends only on the ratio $\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})$. [^7] It is straight-forward to numerically compute the supremum on the right-hand side of the display in 4.4, since the objective function is single-peaked, being the product of an increasing function and a decreasing function. Under correct specification, the probability of the joint event is equal to $a(1-a)$ ($=9\%$ when $a=10\%$). Yet we see that the joint probability can exceed 46% when the asymptotic standard deviation of the VAR estimator is less than half that of the LP estimator. As $\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})\to 0$, the worst-case joint probability approaches $1-a$ [^8] The argument is as follows. On the one hand, the supremum of the function on the right-hand side of the display in 4.4 exceeds the function evaluated at $b=(\operatorname{aVar}(\hat{\beta}_{h})/\operatorname{aVar}(\hat{\delta}_{h})-1)^{1/4}$, and it is easy to see that this function value converges to $1-a$. On the other hand, the first factor in the function is bounded above by 1, and the second factor is bounded above by $1-r(0;z_{1-a/2})=1-a$. We thus again see that statistical tests fail to guard researchers adequately against the potential for severe VAR coverage distortions.
@@ 4.4 Bias-aware inference
To fix the under-coverage of the conventional VAR confidence interval, we can adjust the critical value upward to compensate for the bias, as suggested in a general setting by Armstrong and Kolesár (2021). Suppose again that we restrict the misspecification $\alpha(L)$ to satisfy $\|\alpha(L)\|\leq M$. Then we define the bias-aware VAR confidence interval
@@
where the bias-aware critical value $\operatorname{cv}_{1-a}(b)$ is given by the number such that $r(b;\operatorname{cv}_{1-a}(b))=a$, and $r(\cdot,\cdot)$ is defined in 3.2. By construction, this bias-aware confidence interval has correct (but potentially conservative) asymptotic coverage.
@@
A very tight bound $M$ on the signal-to-noise ratio is required for the bias-aware VAR interval to be shorter than the LP interval. The ratio of the length of the bias-aware interval to that of the conventional LP interval is a function only of $M$ and the asymptotic relative precision of LP and VAR. Figure 4.3 plots the relative interval length as a function of the relative asymptotic standard deviation, for significance level $a=10\%$ and for different misspecification bounds $M$. The figure shows that $M$ has to be quite small—apparently below 1—for the bias-aware VAR length to dominate the LP length regardless of the DGP and horizon. Even for $M=1.5$, bias-aware VAR is at best only moderately shorter than LP (and sometimes longer). For values of $M$ above 2 (corresponding to a noise-to-signal ratio above $4/T$), bias-aware VAR is dominated by LP.
@@
In Section A.2 we show that the conventional LP confidence interval is at most slightly wider than a more efficient bias-aware confidence interval centered at the model averaging estimator $\hat{\theta}_{h}(\omega)=\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$ introduced in 4.2 above. Even if the weight $\omega$ is chosen to optimize confidence interval length, the gains relative to the LP interval are very small when $M\geq 2$ (corresponding to a noise-to-signal ratio above $4/T$).

### v2

@@ 3.2 Hausman misspecification test
Considering separately the three cases $\zeta\in(1/4,1/2)$, $\zeta=1/2$, and $\zeta>1/2$, the result is an immediate consequence of Propositions 3.1 and 3.2 as well as Corollary A.2 in Section A.3. ∎
As claimed previously, the Hausman test is consistent against MA misspecification of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$, except in the knife-edge case where $\operatorname{aBias}(\hat{\delta}_{h})=0$. When $\zeta=1/2$ and $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$, the asymptotic rejection probability is strictly between the significance level $a$ and 1. In Section 4 we will use the Hausman test to quantify the difficulty of detecting especially pernicious types of model misspecification.
3.3 The role of lag length
One simple way to remove the asymptotic bias of the VAR estimator is to control for sufficiently many lags, since in this case the estimator is asymptotically equivalent with the LP estimator. The larger the impulse horizon of interest, the more lags are required for bias reduction. See Plagborg-Møller and Wolf (2021) and Xu (2023) for related results in models without explicit MA misspecification.
@@ 3.3 The role of lag length
We show in Section 4 that controlling for so many lags that LP and VAR are asymptotically equivalent is in fact the only way to guarantee that the asymptotic bias of the VAR estimator is zero.
@@ 4 Some unpleasant VARithmetic
To show that the fragility of VARs is likely to matter in practice, we now investigate the worst-case properties of VAR procedures under a tight constraint on the amount of misspecification. We prove that there is no free lunch: the conventional VAR confidence interval is robust to misspecification if, and only if, the LP and VAR intervals coincide asymptotically. VARs with short-to-moderate lag lengths instead suffer from severe coverage distortions for small amounts of misspecification that is hard to rule out economically or statistically. Beyond increasing the lag length, an alternative strategy to fix VAR undercoverage is to use a larger bias-aware critical value; however, we show that the resulting confidence interval is usually wider than the LP interval. Finally, we generalize all our results to the case of joint inference on multiple impulse responses.
@@
4.1 Worst-case bias and mean-squared error
Building towards our main results on VAR coverage distortions, we begin by deriving the worst-case bias and mean-squared error of the VAR estimator.
Misspecification bound.
@@ 4.1 Worst-case bias and mean-squared error
Suppose we are willing to impose a priori that the noise-to-signal ratio is at most $M^{2}/T$ for some constant $M\in(0,\infty)$. For small $M^{2}/T$, this roughly means that a fraction $M^{2}/T$ of the variance of the model’s error term is due to the misspecification. This corresponds to restricting the parameter space for $\alpha(L)$ to all absolutely summable lag polynomials that satisfy $\|\alpha(L)\|\leq M$. In the following we will consider the worst-case properties of the VAR estimator over this parameter space, treating the other (consistently estimable) parameters $(A,H,D)$ as fixed.
Worst-case bias.
@@
The claim is a special case of Proposition 4.2 below. ∎
Under our bound $M^{2}/T$ on the noise-to-signal ratio, the worst-case (scaled) VAR bias is a simple function of $M$ and of the relative asymptotic precision $\operatorname{aVar}(\hat{\beta}_{h})/\operatorname{aVar}(\hat{\delta}_{h})$ of the VAR estimator vs. LP. These two quantities are “sufficient statistics” for the worst-case bias regardless of the number $n$ of variables in the VAR, the lag length $p$, the specific VAR parameters $(A,H,D)$, and the horizon $h$. Hence, our subsequent analysis of the worst-case properties of VAR procedures depends only on $M$ and on the relative precision, allowing us to concisely present analytical results that cover a wide range of local-to-SVAR models without having to resort to simulations that inevitably only cover a finite number of DGPs.
Proposition 4.1 shows that, in terms of bias, there is “no free lunch” for VAR estimation: the worst-case VAR bias is small precisely when the VAR estimator has nearly the same variance as LP. While the worst-case bias can be reduced by increasing the VAR estimation lag length $p$, the proposition shows that this can only happen at the expense of increasing the variance. If we include so many lags that the worst-case bias is zero (cf. Corollary 3.2), then the VAR estimator must necessarily be asymptotically equivalent with LP.
Worst-case mean squared error.
For future reference we briefly discuss how the worst-case mean squared error (MSE) of the VAR estimator depends on the imposed bound on misspecification. Based on Propositions 3.1 and 3.2 as well as Corollary A.2, we define the asymptotic MSE of the VAR and LP estimators as follows:
@@
See Section B.5. ∎
In words, the worst-case MSE regret of VAR relative to LP is proportional to the variance reduction of VAR relative to LP, with a proportionality constant of $M^{2}-1$. If $M>1$ (corresponding to a noise-to-signal ratio greater than $1/T$), the worst-case MSE of VAR thus strictly exceeds the MSE of LP. From here it is also straightforward to recover the minimax optimal way to average LP and VAR estimates.
@@ 4.2 Worst-case coverage
We now turn to our main area of interest: the worst-case asymptotic coverage of the conventional VAR confidence interval under our bound on the amount of misspecification. This turns out to take a very simple form.
@@
This is an immediate consequence of Corollaries 3.1 and 4.1. ∎
Based on this corollary, Figure 4.1 provides a complete characterization of the robustness-efficiency trade-off for VAR confidence intervals. It plots the worst-case coverage probability as a function of the ratio of standard errors for VAR and LP, given significance level $a=10\%$ and different values of $M$. The shaded area depicts an empirically relevant range of standard error ratios obtained in four empirical applications from Ramey (2016). [^6] We replicate Ramey’s identification schemes for monetary, tax, government spending, and technology shocks. The shaded area shows the 10th to 90th percentiles of standard error ratios at horizons exceeding 1 year. See LABEL:app:emp_se for details. We see that, even for $M=1$ (corresponding to a noise-to-signal ratio of $1/T$), the worst-case coverage probability is below 48% whenever the asymptotic standard deviation of the VAR estimator is less than half that of LP—a value that is typical in applied work. Further, at the bottom end of the empirically relevant range, the worst-case coverage probability is essentially zero as soon as $M\geq 1$. It is only at the very right side of the figure—when the VAR includes enough lags to remove nearly all bias, thus increasing the standard error almost to that of LP—that the VAR confidence interval has coverage close to the nominal level.
Figure 4.1: Worst-case asymptotic coverage probability of the conventional 90% VAR confidence interval. Horizontal axis: relative asymptotic standard deviation of LP vs. VAR. Different lines: different bounds $M$ on $\|\alpha(L)\|$. Shaded area: empirical 10th–90th percentile range of relative standard errors based on Ramey (2016), see LABEL:app:emp_se. The solid horizontal line marks the nominal coverage probability $1-a=90\%$.
The potential for VAR undercoverage documented here may not be so concerning if the worst-case misspecification can be ruled out on economic theory grounds, or if it is easily detectable statistically. We now argue that neither appears to be the case.
@@
The shape and magnitude of the least favorable misspecification is difficult to rule out generally based on economic theory. The least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha_{\ell,h,M}^{\dagger}L^{\ell}$ for VAR coverage is the same as the least favorable one for bias (i.e., the $\alpha(L)$ that achieves the maximum in Proposition 4.1). Since $\operatorname{aBias}(\hat{\delta}_{h})$ is linear in $\alpha(L)$, the least favorable choice given the constraint $\|\alpha(L)\|\leq M$ follows easily from the Cauchy-Schwarz inequality (see the proof of Proposition 4.2 below):
$\alpha_{\ell,h,M}^{\dagger}\propto D^{1/2}H^{\prime}\Psi_{h}^{\prime}S^{-1}A^{\ell-1}HD^{1/2}-\mathbbm{1}(\ell\leq h)\sigma_{j^{*}}^{-1}D^{1/2}H^{\prime}(A^{\prime})^{h-\ell}e_{i^{*},n}e_{j^{*},m}^{\prime},\quad\ell\geq 1,$ (4.1)
@@
We note two main properties of the least favorable misspecification. First, the magnitude of the MA coefficients $\alpha_{\ell,h,M}^{\dagger}$ decays exponentially as $\ell\to\infty$. In other words, not only is the overall magnitude of the least favorable model misspecification small (as imposed in the noise-to-signal bound), the MA coefficients at long lags are in fact particularly small. Second, numerical examples shown in Section A.1 suggest that the MA coefficients tend to be largest in magnitude at horizon $h$, displaying either a hump-shaped pattern as a function of $\ell$—consistent with economic theories of adjustment costs or learning—or a single zig-zag pattern—consistent with theories of overshooting or lumpy adjustment. We thus view MA dynamics of the worst-case form as empirically and theoretically relevant types of misspecification. [^7] However, the least favorable MA polynomial derived above need not be of interest to researchers who trust that some equations in their SVAR specification are exactly correctly specified, as this imposes the additional restrictions that some linear combinations of the rows of the MA polynomial $\alpha(L)$ equal zero.
@@
Rather than committing a priori to a parameter space for $\alpha(L)$ through choice of $M$, we can also ask a different question: across all possible types and magnitudes of misspecification, what is the worst-case probability that the conventional VAR confidence interval fails to cover the true impulse response, yet we fail to reject correct specification of the VAR model?
@@
See Section B.7. ∎
@@
Figure 4.2 plots this worst-case probability for a significance level of $a=10\%$, which by Corollary 4.4 depends only on the ratio $\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})$. Under correct specification, the probability of the joint event is equal to $a(1-a)$ ($=9\%$ when $a=10\%$). With misspecification, the joint probability instead exceeds 46% when the asymptotic standard deviation of the VAR estimator is less than half that of the LP estimator. As $\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})\to 0$, the worst-case joint probability approaches $1-a$. We thus again see that statistical tests may fail to warn against the potential for severe VAR coverage distortions.
@@ 4.3 Bias-aware inference
Rather than removing bias by increasing the lag length (thus ensuring equivalence with LP), an alternative way to fix the undercoverage of the conventional VAR confidence interval is to adjust the critical value upward to compensate for the bias, as suggested in a general setting by Armstrong and Kolesár (2021). Suppose again that we restrict the misspecification $\alpha(L)$ to satisfy $\|\alpha(L)\|\leq M$. Then we define the bias-aware VAR confidence interval
@@
where the bias-aware critical value $\operatorname{cv}_{1-a}(b)$ is given by the number such that $r(b;\operatorname{cv}_{1-a}(b))=a$, and $r(\cdot,\cdot)$ is defined in Corollary 3.1. By construction, this bias-aware confidence interval has correct (but potentially conservative) asymptotic coverage.
@@
It turns out, however, that a very tight bound $M$ on the signal-to-noise ratio is required for the bias-aware VAR interval to be shorter than the LP interval. Figure 4.3 plots the relative interval length as a function of the relative asymptotic standard deviation of VAR and LP, for a significance level of $a=10\%$ and for different misspecification bounds $M$. The figure shows that $M$ has to be quite small—apparently below 1—for the bias-aware VAR length to dominate the LP length regardless of the DGP and horizon. Even for $M=1.5$, bias-aware VAR is at best only moderately shorter than LP. Finally, for values of $M$ above 2 (corresponding to a noise-to-signal ratio above $4/T$), bias-aware VAR is dominated by LP.
@@
In Section A.2 we furthermore show that the conventional LP confidence interval is at worst slightly wider than a more efficient bias-aware confidence interval centered at the model averaging estimator $\hat{\theta}_{h}(\omega)=\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$, introduced in Corollary 4.2 above. Even if the weight $\omega$ is chosen to optimize confidence interval length, the gains relative to the LP interval are very small when $M\geq 2$ (corresponding to a noise-to-signal ratio above $4/T$).